# Analysis of a Programmable Quantum Annealer as a Random Number Generator

### 19 Dec 2023 v2

@@ 1 Introduction
One of the types of programmable quantum computers that have become available to test, typically as cloud computing resources, are D-Wave quantum annealers. Quantum annealing is a specialized type of quantum computation that aims to sample the optimal solution of a combinatorial optimization problem, and is typically implemented using the transverse driving Hamiltonian [6, 7, 8, 9, 10, 11]. D-Wave quantum annealers are physically implemented using programmable superconducting flux qubits [12, 13, 14, 15, 16, 17, 18]. Quantum annealers, and more generally quantum computers, are potentially interesting as secure entropy sources for generating random numbers because of the inherent stochasticity of quantum states - there is not a deterministic mechanism to compute what the measured state will be of an arbitrary quantum state. For this reason, quantum computers, and more generally physical sources of measurements of quantum information, are True Random Number Generators (TRNGs) (or QRNGs) [19, 20, 4]. Importantly, there exist current technologies which are secure, high bit-rate, QRNGs [21, 22].
@@
The primary reason that modern quantum annealers are not perfect random number generators is because there are a large number of sources of error and bias in the computation, for example the spin bath polarization effect [23, 24] can cause sequential anneal-readout cycles to have self correlations (in time), and programmed coefficients (even if they are 00) have slightly different effective weights on the hardware [25]. Furthermore, it has been shown that modern D-Wave quantum annealers have a measurable performance change over time [26, 27]. There have also been cross-qubit correlations observed on a D-Wave 2000Q chip [28, 29]. There have been studies which aim to reduce biases and noise present in minor-embedded QA computations, which in the case of reducing biases in the constraint of the graph partitioning problem results in effectively attempting to create an unbiased Quantum Annealing random number sampler, see ref. [30]. Interestingly, quantum annealing (even in an ideal computation, with no noise), does not sample degenerate groundstates (i.e. optimal solutions of the combinatorial optimization problem) uniformly due to the transverse field driving Hamiltonian [31, 32, 33, 34, 35], meaning that using the QA sampling of the groundstates of non-trivial Hamiltonians would not be a good source of unbiased random numbers. Instead, the much simpler case of an all 00 coefficient Ising model is the most direct way to program these devices to produce random numbers (see more details in Section 2.1). D-Wave quantum annealers have been evaluated, on somewhat small problem sizes, for the possibility of utilizing them as TRNGs  [^1] Sometimes the acronym that is used for quantum devices generating random numbers is Quantum Random Number Generator (QRNG) in previous studies [36, 37].
@@ 2 Methods
Section 2.1 details the Quantum Annealing implementation details, and Section 2.2 details the randomness testsuite that is used.
@@ 2.1 Quantum Annealing
The computation performed by D-Wave quantum annealers is described by eq. (1), and eq. (2) describes the discrete optimization Ising model that a user can program to be sampled by the quantum annealer (the quadratic coefficients are subject to the constraint of the native connectivity of the quantum annealing hardware). Eq. (2) is just a slight re-formulation of the Ising model defined in the second summation term of eq. (1). The goal of the quantum annealer is to find a minimum variable assignment vector $z$ given the objective function Eq.(2). The variable states can be either $\{0,1\}^{n}$, in which case the combinatorial optimization problem is called a Quadratic Unconstrained Binary Optimization (QUBO) problem, or the variables can be spins $\{+1,-1\}^{n}$, in which case the combinatorial optimization problem is called an Ising model.
@@
The D-Wave quantum annealer that is used to generate random bits is DW_2000Q_LANL, the Chimera hardware graph for this device is shown in Figure 1. The simplest way to generate random bits using a quantum annealer is simply to set the user programmed coefficients for all linear terms (e.g. hardware qubits) and quadratic terms (e.g. hardware couplers) to 00, meaning that only the transverse field terms are present in the computation (ideally), which means that the qubits are in a uniform superposition, while the computation is coherent, during the anneal. There are certainly more complicated ways that could be utilized with the goal of extracting good random bits, such as random circuit sampling on gate model devices [51] or by tuning devices biases to improve sampling of balanced partitions for the graph partitioning problem [30], however in general RNG’s need to be as fast as possible and therefore minimizing the complexity of the computation is likely a good motivation to aim for. Explicitly, the Ising model (variable states are $\in\{+1,-1\}$) that we will sample is given in eq. (3).
@@
The main D-Wave parameters that will be varied are the annealing time, which will use $1$, $10$, $100$, and $2000$ microseconds. These annealing times span the range of allowed annealing times on the DW_2000Q_LANL chip; $1$ microsecond is the smallest annealing time and $2000$ is the longest available annealing time. The other parameter that will be tested is turning on server side classical post processing which aims to improve sampling (although in this specific case, the sampling is being done on an all zero coefficient Ising model). In order to turn on this server side post processing, the user facing post process option was set to sampling [52]. The sampling server side post-processing option performs local bit changes on the measurements (before sending the results to the user) with the goal of obtaining a post-processes set of samples that corresponds to a Boltzmann distribution with inverse temperature $\beta$, where $\beta$ is $\beta$ is set to a value near to the inverse temperature corresponding to the raw samples (see ref. [52] for more details). We test turning this server-side post processing option on and off. The reasoning is that we would ideally want to the quantum annealer to be able to produce unbiased random bits without this post processing, however it may be the case that the classical post processing helps reduce bias in the samples at with a small computational overhead, in which case it would be interesting to quantify this. Therefore, in total there will be $8$ datasets, each using a different quantum annealer parameter choice. Each of the datasets will be strictly sequential in time - e.g. the order of the bits will not be changed by some other entropy source. This time series representation of the data is especially important since it has been shown that there are long term trends that can be observed in current D-Wave quantum annealers [26]. Additionally, the exact ordering of the bits within each anneal (e.g. whole-chip readout cycle) is strictly based on the logical qubit indexing within the hardware, which is fixed for all samples, but is arbitrarily set. Each anneal-readout cycle is concatenated with the next anneal-readout cycle that was executed in time - no other source of entropy is present in the data.
In order to mitigate spin bath polarization effect [23] self-sample correlations, all of the data is constructed by sequentially calling the D-Wave backend for a single anneal-readout cycle (i.e. instead of measuring many anneals in a single job). Each job is sampled as an Ising model, meaning that spins are the measured states (although whether the model was specified as QUBO or Ising would in principle have no impact on the results). Additionally, although the bits are all time ordered, because of network interruptions or device power losses there are gaps in the time of the sequential random bit sampling.
@@
For all $8$ datasets, the programming_thermalization and readout_thermalization are set to 00 microseconds in order to remove any thermalization effects (beyond thermalization that occurs after the qubits lose coherence; the qubit coherence times are estimated to be on the order of 10’s of nanoseconds [53]). All other parameters are set to default. The motivation for evaluating these different parameter choices is the following.
@@
    Although in general increasing annealing times on D-Wave quantum annealers results in better sampling success rate of combinatorial optimization problems, these long annealing times are well out of the qubit coherence times of current D-Wave quantum annealers [54, 53]. Therefore, the longer annealing times are using thermalization to marginally improve the sampling success probability [55]. However, in this case there is not a combinatorial optimization problem being sampled - therefore in principle it may be the case that longer anneal times accumulate more errors in the computation, in particular more biases in the random sampling computation we aim to perform. Therefore, it may be advantageous to sample using the shortest annealing times available on the hardware – whether this is true or not is the aim of the varying annealing times. Furthermore, sampling rate for random numbers is extremely important – faster random bit sampling is more useful than slower sampling. Therefore, if the smaller annealing times produce high entropy random bits this would be better than using longer annealing times.
@@ 2.2 Testing for randomness
The randomness test that will be applied to the data are all of the tests from SP800-22 Rev 1a by National Institute of Standards and Technology [56], titled A Statistical Test Suite for Random and Pseudorandom Number Generators for Cryptographic Applications. This testsuite contains 15 randomness tests, two of which contain several sub-tests. In total each of the $8$ datasets will be tested against $38$ randomness tests, each giving a p-value output. For the purposes of maintaining consistency, and using the original NIST SP 800-22 test definitions [56], a computed P-value which is $\geq 0.01$ would accept the sequence as being random, and otherwise we would consider it to be non-random. This p-value threshold criteria is applied to all of the randomness tests. In the tests where there are multiple computed P-values, such as cumulative sums where there is a forward and backward mode of operation, all P-values must be greater than or equal to $0.01$ in order to be considered random. The serial test also outputs two P-values.
@@ 4 Discussion and Conclusion
A potentially interesting question is to numerically examine how ideal quantum annealing, with no noise or environment interaction, would sample the same Ising model with 00 Ising coefficients (e.g. only the transverse field Hamiltonian).
Another potentially interesting analysis on this existing data from DW_2000Q_LANL would be to determine if there are strong cross-qubit correlations on the chip. If such correlations exist, then this would indicate cross-talk errors.

### 02 Feb 2024 v3

@@ 1 Introduction
One of the types of programmable quantum computers that have become available to test, typically as cloud computing resources, are D-Wave quantum annealers. Quantum annealing is a specialized type of quantum computation that aims to sample the optimal solution(s) of a combinatorial optimization problem, ideally using adiabatic evolution [6, 7, 8, 9, 10]. Quantum annealing hardware is typically implemented using the transverse driving Hamiltonian where the system is initialized in the groundstate of the Transverse-field Hamiltonian [11, 8, 9, 10, 7, 12]. D-Wave quantum annealers are physically implemented using programmable superconducting flux qubits [13, 14, 15, 16, 17, 18, 19]. Quantum annealers, and more generally quantum computers, are potentially interesting as secure entropy sources for generating random numbers because of the inherent stochasticity of measuring quantum states - there is not a deterministic mechanism to compute what the measured state will be of an arbitrary quantum state. For this reason, quantum computers, and more generally physical sources of measurements of quantum information, are True Random Number Generators (TRNGs) (or QRNGs) [20, 21, 4]. Importantly, there exist current technologies which are secure, high bit-rate, QRNGs [22, 23].
@@
The primary reason that modern quantum annealers are not perfect random number generators is because there are a large number of sources of error and bias in the computation, for example the spin bath polarization effect [24, 25] can cause sequential anneal-readout cycles to have self correlations (in time), and programmed coefficients (even if they are 00) have slightly different effective weights on the hardware [26]. Furthermore, it has been shown that modern D-Wave quantum annealers have a measurable performance change over time [27, 28]. There have also been cross-qubit correlations observed on a D-Wave 2000Q chip [29, 30]. There have been studies which aim to reduce biases and noise present in minor-embedded QA computations, which in the case of reducing biases in the constraint of the graph partitioning problem results in effectively attempting to create an unbiased Quantum Annealing random number sampler, see ref. [31]. Interestingly, quantum annealing (even in an ideal computation, with no noise), does not in general sample degenerate groundstates (i.e. optimal solutions of the combinatorial optimization problem) uniformly due to the transverse field driving Hamiltonian [32, 33, 34, 35, 36], meaning that using the QA sampling of the groundstates of non-trivial Hamiltonians would not be a good source of unbiased random numbers. Instead, the much simpler case of an all 00 coefficient Ising model is the most direct way to program these devices to produce random numbers (see more details in Section 2.1). D-Wave quantum annealers have been evaluated, on somewhat small problem sizes, for the possibility of utilizing them as TRNGs  [^1] Sometimes the acronym that is used for quantum devices generating random numbers is Quantum Random Number Generator (QRNG) in previous studies [37, 38].
@@
All data from this study are publicly available as a Zenodo dataset [53].
@@ 2 Methods
Section 2.1 details the Quantum Annealing implementation details, and Section 2.2 details the randomness testsuites that are used.
@@ 2.1 Quantum Annealing
The computation performed by D-Wave quantum annealers is described by eq. (1), and eq. (2) describes the discrete optimization Ising model that a user can program to be sampled by the quantum annealer (the quadratic coefficients are subject to the constraint of the native connectivity of the quantum annealing hardware). The functions $A(s)$ and $B(s)$ define the Transverse field driving Hamiltonian strength and the programmed Ising model Hamiltonian, respectively, parameterized by the variable $s$. In standard quantum annealing, which is the setting used in this study, $s$ defines a linear schedule as a function of anneal time, and the strengths of $A(s)$ and $B(s)$ at each $s$ step are system defined quantities. At the beginning of the anneal the $A(s)$ term dominates, and then over the course of the anneal $A(s)$ is reduced in strength and $B(s)$ is increased in strength. Eq. (2) is a slight re-formulation of the Ising model defined in the second summation term of eq. (1). The goal of the quantum annealer is to find a minimum variable assignment vector $z$ given the objective function eq. (2). The variable states can be either $\{0,1\}^{n}$, in which case the combinatorial optimization problem is a Quadratic Unconstrained Binary Optimization (QUBO) problem, or the variables can be spins $\{+1,-1\}^{n}$, in which case the combinatorial optimization problem is an Ising model.
@@
The D-Wave quantum annealer that is used to generate random bits is DW_2000Q_LANL, the Chimera hardware graph for this device is shown in Figure 1. The simplest way to generate random bits using a quantum annealer is simply to set the user programmed coefficients for all linear terms (e.g. hardware qubits) and quadratic terms (e.g. hardware couplers) to 00, meaning that only the transverse field Hamiltonian is present in the computation (ideally), which means that the qubits are in a uniform superposition, while the computation is coherent, during the anneal. There are certainly more complicated ways that could be utilized with the goal of extracting good random bits, such as random circuit sampling on gate model devices [52] or by tuning devices biases to improve sampling of balanced partitions for the graph partitioning problem [31], however in general RNG’s need to be as fast as possible and therefore minimizing the complexity of the computation is likely a good motivation to aim for. Explicitly, the Ising model (variable states are $\in\{+1,-1\}$) that we will sample is given in eq. (3).
@@
The main D-Wave parameters that will be varied are the annealing time, which will use $1$, $10$, $100$, and $2000$ microseconds. These annealing times span the range of allowed annealing times on the DW_2000Q_LANL chip; $1$ microsecond is the smallest annealing time and $2000$ is the longest available annealing time. The other parameter that will be tested is turning on server side classical post processing which aims to improve sampling (although in this specific case, the sampling is being done on an all zero coefficient Ising model). In order to turn on this server side post processing, the user facing post process option was set to sampling [54]. The sampling server side post-processing option performs local bit changes on the measurements (before sending the results to the user) with the goal of obtaining a post-processes set of samples that corresponds to a Boltzmann distribution with inverse temperature $\beta$, where $\beta$ is $\beta$ is set to a value near to the inverse temperature corresponding to the raw samples (see ref. [54] for more details). We test turning this server-side post processing option on and off. The reasoning is that we would ideally want to the quantum annealer to be able to produce unbiased random bits without this post processing, however it may be the case that the classical post processing helps reduce bias in the samples at with a small computational overhead, in which case it would be interesting to quantify this. Therefore, in total there will be $8$ datasets, each using a different quantum annealer parameter choice. Each of the datasets will be strictly sequential in time - e.g. the order of the bits will not be changed by some other entropy source. This time series representation of the data is especially important since it has been shown that there are long term trends that can be observed in current D-Wave quantum annealing processors [27]. Additionally, the exact ordering of the bits within each anneal (e.g. whole-chip readout cycle) is strictly based on the logical qubit indexing within the hardware, which is fixed for all samples, but is arbitrarily set. Each anneal-readout cycle is concatenated with the next anneal-readout cycle that was executed in time - no other source of entropy is present in the data.
With the goal of mitigating the spin bath polarization effect [24] self-sample correlations, all of the data is constructed by sequentially calling the D-Wave backend for a single anneal-readout cycle (i.e. instead of measuring many anneals in a single job). Each job is sampled as an Ising model, meaning that spins are the measured states (although whether the model was specified as QUBO or Ising would in principle have no impact on the results). Additionally, although the bits are all time ordered, because of network interruptions or device power losses there are gaps in the time of the sequential random bit sampling.
@@
For all $8$ datasets, the programming_thermalization and readout_thermalization are set to 00 microseconds so as to remove any thermalization effects (beyond thermalization that occurs after the qubits lose coherence; the qubit coherence times are estimated to be on the order of 10’s of nanoseconds [55]). All other parameters are set to default. The motivation for evaluating these different parameter choices is the following.
@@
    Although in general increasing annealing times on D-Wave quantum annealers results in better sampling success rate of combinatorial optimization problems, these long annealing times are much longer than the qubit coherence times of current D-Wave quantum annealers [56, 55]. Therefore, the longer annealing times are using thermalization to marginally improve the sampling success probability [57]. However, in this case there is not a combinatorial optimization problem being sampled - therefore in principle it may be the case that longer anneal times accumulate more errors in the computation, in particular more biases in the random sampling computation we aim to perform. Therefore, it may be advantageous to sample using the shortest annealing times available on the hardware – whether this is true or not is the aim of the varying annealing times. Furthermore, sampling rate for random numbers is extremely important – faster random bit sampling is more useful than slower sampling. Therefore, if the smaller annealing times produce high entropy random bits this would be better than using longer annealing times.
@@ 2.2 Testing for randomness
The randomness test that will be applied to the data are all of the tests from SP800-22 Rev 1a by National Institute of Standards and Technology [58], titled A Statistical Test Suite for Random and Pseudorandom Number Generators for Cryptographic Applications. This testsuite contains 15 randomness tests, two of which contain several sub-tests. In total each of the $8$ datasets will be tested against $38$ randomness tests, each giving a p-value output. For the purposes of maintaining consistency, and using the original NIST SP 800-22 test definitions [58], a computed P-value which is $\geq 0.01$ would accept the sequence as being random, and otherwise we would consider it to be non-random. This p-value threshold criteria is applied to all of the randomness tests. In the tests where there are multiple computed P-values, such as cumulative sums where there is a forward and backward mode of operation, all P-values must be greater than or equal to $0.01$ to be considered random. The serial test also outputs two P-values.
@@ 4 Discussion and Conclusion
A potentially interesting analysis on this existing data from DW_2000Q_LANL would be to determine if there are strong cross-qubit correlations on the chip. If such correlations exist, then this could indicate cross-talk errors from the control system.